# FunDiff: Diffusion Models over Function Spaces for Physics-Informed Generative Modeling

### v1

@@ Abstract
Recent advances in generative modeling—particularly diffusion models and flow matching—have achieved remarkable success in synthesizing discrete data such as images and videos. However, adapting these models to physical applications remains challenging, as the quantities of interest are continuous functions governed by complex physical laws. Here, we introduce FunDiff, a novel framework for generative modeling in function spaces. FunDiff combines a latent diffusion process with a function autoencoder architecture to handle input functions with varying discretizations, generate continuous functions evaluable at arbitrary locations, and seamlessly incorporate physical priors. These priors are enforced through architectural constraints or physics-informed loss functions, ensuring that generated samples satisfy fundamental physical laws. We theoretically establish minimax optimality guarantees for density estimation in function spaces, showing that diffusion-based estimators achieve optimal convergence rates under suitable regularity conditions. We demonstrate the practical effectiveness of FunDiff across diverse applications in fluid dynamics and solid mechanics. Empirical results show that our method generates physically consistent samples with high fidelity to the target distribution and exhibits robustness to noisy and low-resolution data. Code and datasets are publicly available at https://github.com/sifanexisted/fundiff.
Keywords generative modeling  $\cdot$ diffusion models  $\cdot$ partial differential equations  $\cdot$ physics-informed machine learning
1 Introduction
@@
Generative modeling, a cornerstone of modern machine learning, focuses on learning and sampling complex data distributions $P(x)$ by constructing models that can generate new instances consistent with the learned distribution [1, 2]. Among recent advances in this domain, diffusion models and flow matching have emerged as powerful frameworks that achieve state-of-the-art performance. Diffusion models [3, 4] learn to reverse a forward noising process by gradually denoising data through a Markov chain, which iteratively transforms random noise into high-quality samples. In parallel, flow matching [5, 6] provides an alternative approach by learning continuous normalizing flows via optimal transport theory, directly modeling probability paths that transform a tractable prior distribution into the target data distribution.
@@
These methodologies have demonstrated exceptional capabilities across diverse applications, from high-resolution image synthesis [7, 8, 9] and audio generation [10] to protein structure prediction [11] and molecular design [12]. However, while these advances have primarily focused on discrete data structures such as images, text, and molecular graphs, many scientific applications, particularly in physics, require generating continuous function data where each sample represents an entire function or field. For instance, generating velocity fields in turbulent fluid dynamics, simulating deformation patterns in complex materials under external loadings, and forecasting temperature variations in climate models all require handling functions defined over continuous spatial and temporal domains [13]. These problems inherently involve infinite-dimensional function spaces, a challenge that recent advances in operator learning frameworks have begun to address effectively [14, 15, 16, 17, 18]. This presents unique challenges:
@@
How can we extend generative modeling to function-valued data, while preserving important structural or physical properties inherent in scientific problems?
@@
Recent works have attempted to extend generative models to function data, either by adapting existing algorithms to work with discretized function data in $\mathbb{R}^{n}$ or by transforming input functions into a finite-dimensional latent space [19, 20, 21, 22, 23, 24]. Dupont et al. [25] pioneered this direction by treating data points as continuous functions and learning distributions over them using adversarial training. Multiple theoretical frameworks have emerged for function diffusion models, including Gaussian measures on Hilbert spaces [26], dimension-independent bounds for infinite-dimensional models [27], generalized diffusion processes in function spaces [28], and denoising diffusion operators with rigorous Gaussian processes and Langevin dynamics [29]. Although these approaches have shown promise for fixed discretizations, they face challenges in scaling to higher resolutions and handling variable discretizations. Moreover, they lack a principled foundation for extending generative modeling to function spaces, where each sample represents a continuous object rather than a finite-dimensional vector. As a result, these approaches do not scale naturally to problems where the underlying data lies in an infinite-dimensional space, such as those encountered in scientific computing, partial differential equation (PDE) modeling, or operator learning [15, 13, 30].
@@
In parallel, diffusion models have emerged as powerful tools for tackling inverse problems and modeling dynamical systems governed by PDEs. Recent works have developed methods for reconstructing spatiotemporal fields from sparse observations and partial supervision [31, 32, 33]. To address resolution limitations, several approaches have integrated neural operators with diffusion models [34, 35, 36], demonstrating improved performance in turbulence modeling and multi-scale phenomena capture. Furthermore, various strategies have attempted to incorporate physical constraints directly into diffusion processes [37, 38, 39, 40, 32], either by modifying the training objective or by projecting generated samples onto constraint manifolds during posterior sampling. However, many physical constraints are inherently formulated in continuous space but must be discretized during training due to the discrete nature of these models. This discretization introduces numerical artifacts, often resulting in inaccurate or ill-defined PDE residuals, especially in inverse problems where key variables are partially or entirely unobserved. These challenges underscore a fundamental gap between current data-driven generative models and the objective of producing physically consistent, function-valued outputs.
To address these challenges, we propose FunDiff, a novel framework for physics-informed generative modeling over function spaces. Our contributions are as follows.
  • We introduce a unified architecture that combines latent diffusion processes with function autoencoders, enabling the generation of continuous functions from inputs with varying discretizations. This design allows for evaluation at arbitrary spatial locations while seamlessly incorporating physical priors through the function space representation.
  • We establish theoretical guarantees for our approach by proving that diffusion-based models attain minimax-optimal rates for density estimation in infinite-dimensional spaces under mild smoothness assumptions.
  • We demonstrate the practical effectiveness of FunDiff across diverse applications in fluid dynamics and solid mechanics. Our experiments show that the method generates physically consistent samples with high fidelity to target distributions while maintaining robustness to noisy and low-resolution input data.
Taken all together, our framework bridges the gap between theoretical guarantees in function spaces and practical neural network algorithms, establishing a principled approach to physics-informed generative modeling.
The remaining of our paper is structured as follows. In Section 2, we provide an overview of diffusion models, highlighting key concepts and recent advances. Section 3 presents our theoretical results on generative models in function spaces, establishing the mathematical foundation of our work. Building on this theory, we introduce FunDiff, a practical framework for generating functions while incorporating physical priors. In Section 4, we validate our approach with a series of examples, demonstrating that our method is capable of producing physically consistent and high-fidelity continuous functions. Finally, in Section 5, we summarize our findings and discuss potential directions for future research. For notations, please refer to the Appendix A.
2 Preliminaries
Before introducing our generative models in function spaces, we provide some basic knowledge in diffusion models as well as its minimax optimality.
2.1 Diffusion models
We begin by reviewing foundational concepts of diffusion models in both discrete and continuous settings. Diffusion models constitute a class of generative frameworks that approximate complex data distributions through a two-phase process: a forward noising process that incrementally corrupts the data, followed by a backward denoising process that reconstructs data from noise.
$q({\mathbf{x}}_{1:N}\mid{\mathbf{x}}_{0})=\prod_{k=1}^{N}q({\mathbf{x}}_{k}\mid{\mathbf{x}}_{k-1}),\quad\text{with}\quad q({\mathbf{x}}_{k}\mid{\mathbf{x}}_{k-1})=\mathcal{N}(a_{k}{\mathbf{x}}_{k-1},b_{k}^{2}\bm{I}),$
$\frac{\mathrm{d}{\mathbf{x}}}{\mathrm{d}t}=\mathbf{v}({\mathbf{x}},t),$
2.2 Density estimation and the minimaxity of diffusion models
As a core problem in statistics and machine learning, the purpose of density estimation is to solve the following problem: given $n$ i.i.d. data points $\{x_{i}\}_{i=1}^{n}$ sampled from an unknown distribution $p^{*}$, we want to construct an approximate distribution $\hat{p}$ based on the $n$ given data points. Density estimation arises in diverse applications, from anomaly detection and clustering to modern generative modeling, where having an accurate handle on the underlying distribution is crucial. One broad class of solutions is parametric methods, which assume that the ground truth distribution $p^{*}$ lies within a known family of distributions $\{p_{\theta}:~{}\theta\in\Theta\}$. The density estimation problem then reduces to finding the optimal parametric vector $\hat{\theta}$ via log-likelihood maximization:
$\hat{p}_{h}(x)=\frac{1}{nh^{d}}\sum_{i=1}^{n}K\left(\frac{x-x_{i}}{h}\right).$
Recently, diffusion models have emerged as powerful generative approaches that approximate the data distribution by reversing a forward noise-adding process. Although widely recognized for their empirical success, these models also admit a theoretical interpretation under the lens of minimax optimality for density estimation. In fact, under suitable assumptions on the function space that characterizes the ground truth distribution $p^{*}$, one can show that diffusion-based estimators achieve minimax rates for distribution (or score) recovery. Here we outline the theoretical underpinnings of this minimax perspective by first presenting a lower bound result (Theorem 2.1) that establishes a fundamental limit for density estimation in the space of all continuous functions $\mathcal{C}({\mathbb{R}}^{d})$, and then provide an upper bound result (Theorem 2.2) that shows how diffusion-based methods (and related estimators) can achieve this rate, thus elucidating the minimax guarantees behind modern generative modeling frameworks.
Theorem 2.1 (Lower Bound, Theorem 3 of Ref. [45]).
  • (i)
  • (ii)
Theorem 2.2 (Upper Bound, Section 3.3 of Ref. [46]).
Let $\alpha\geqslant 0$ denote the smoothness parameter, $d\geqslant 1$ the dimension, and $c>0$ a positive constant. We still consider the class of densities $\mathcal{F}\subseteq\mathcal{C}^{\alpha}({\mathbb{R}}^{d})$ defined in Theorem 2.1. By solving the backward diffusion SDE
${\mathrm{d}}Y_{t}=\hat{s}(Y_{t},T-t){\mathrm{d}}t+{\mathrm{d}}W_{t}$
3 Generative models over function spaces
@@
In this section, we introduce a framework for generative modeling in function spaces. Since these spaces are inherently infinite-dimensional and thus challenging to handle directly, we first reduce the problem to a finite-dimensional setting. By embedding functions into a finite-dimensional vector space, the task of generating new functions can be reformulated as generating new vectors, an extensively studied and well-understood problem.
3.1 Overview
@@
where the expectation is taken over the randomness of the samples, and $\varepsilon(d,n)$ aligns with the known minimax estimation rates [47, 46]. For the continuous density class $\mathcal{C}({\mathbb{R}}^{d}):=\{\mu:{\mathbb{R}}^{d}\rightarrow[0,\infty)~{}:~{}\mu~{}\text{is continuous}\}$, we have a minimax density estimation rate
@@
In order to measure the distribution loss, we apply an analogy of the Wasserstein distance $W_{1}(P,\widehat{P})$. Notice that for distributions over the function space, the Wasserstein distance is defined similarly as
@@
In the next part, we propose an upper bound of $W_{1}(P,\widehat{P})$ and explain all its factors.
3.2 Density estimation loss decomposition
@@
$W_{1}(P,\widehat{P})\leqslant{\mathbb{E}}_{f\sim P}\|f-\mathcal{D}\circ\mathcal{E}(f)\|_{2}+\mathrm{Lip}(\mathcal{D})\cdot W_{1}(p,\widehat{p}).$ (3.1)
@@
This quantity measures the robustness of $\mathcal{D}$ with respect to perturbations in the input vector space. The third term, $W_{1}(p,\widehat{p})$, reflects the accuracy of the density estimation for vector space, highlighting the performance of the oracle $\mathcal{O}(D,n)$ in estimating the true distribution $p$. The dimension of the feature space $D$ must be chosen judiciously to strike a balance between ensuring accurate autoencoder reconstruction and mitigating the effects of the curse of dimensionality for the vector-space generative oracle. The proof of Eq. (3.1) can be seen in Appendix B.1. Next, we propose some simple examples to help us understand the theoretical framework.
3.2.1 Illustrative example 1: Reproducing kernel Hilbert space
For a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, denote its eigenvalues as $\mu_{1}\geqslant\mu_{2}\geqslant\ldots$, and their corresponding eigenfunctions as $\psi_{1},\psi_{2},\ldots$. For any function $f\in\mathcal{H}$, its eigen-decomposition has the form $f=\sum_{i=1}^{\infty}\theta_{i}\sqrt{\mu_{i}}\psi_{i}$ and its RKHS norm is defined as $\|f\|_{\mathcal{H}}^{2}=\sum_{i=1}^{\infty}\theta_{i}^{2}$. Define the RKHS unit ball as :=$\mathcal{B}:=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leqslant 1\}$.
@@ 3.2.1 Illustrative example 1: Reproducing kernel Hilbert space
$f^{(k)}=\sum_{i=1}^{\infty}\theta_{i}^{(k)}\cdot\sqrt{\mu_{i}}\psi_{i},~{}~{}~{}(k=1,2,\ldots,n).$
@@
$\mathcal{E}:f\mapsto(\theta_{1},\theta_{2},\ldots,\theta_{D})^{\top},~{}~{}~{}\mathcal{D}:(\theta_{1},\theta_{2},\ldots,\theta_{D})^{\top}\mapsto\sum_{i=1}^{D}\theta_{i}\cdot\sqrt{\mu_{i}}\psi_{i}.$
Hence, the autoencoder $\mathcal{D}\circ\mathcal{E}$ effectively serves as a projection (truncation) onto the subspace spanned by the first $D$ eigenfunctions. Next, in order to upper bound the density estimation loss $W_{1}(P,\widehat{P})$, we introduce two different assumptions corresponding to the polynomial and exponential eigenvalue decay of the RKHS $\mathcal{H}$.
@@
Theorem 3.1 (Density estimation over RKHS space).
Assume $P$ is a distribution supported on the RKHS unit ball :=$\mathcal{B}:=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leqslant 1\}$. By using the generative oracle $\mathcal{O}(D,n)$ and the function autoencoder (FAE) framework introduced above, the density estimation loss over RKHS space can be bounded as follows:
@@
The proof is provided in Appendix B.2. ∎
3.2.2 Illustrative example 2: Barron space
@@
Definition 3.2 (Path norm [49]).
@@
Definition 3.3 (Spectral norm).
@@
$f({\mathbf{x}})=\int_{{\mathbb{R}}^{d}}e^{i\langle{\mathbf{x}},\omega\rangle}\hat{F}(\omega)\,{\mathrm{d}}\omega,~{}~{}\forall{\mathbf{x}}\in\Omega.$
@@
Lemma 3.4.
For any $f\in\mathcal{F}_{s}$, there exists a two-layer network $f({\mathbf{x}};\tilde{\theta})$ of width $m$ such that:
@@
$\|{\mathbf{b}}_{k}\|_{1}=1,|c_{k}|\leqslant 1,~{}~{}\text{and}~{}~{}|a_{k}|\leqslant\frac{2\gamma(f)}{m}\leqslant\frac{2}{m}~{}~{}~{}~{}\text{holds for~{}~{}}\forall k\in[m].$
In the context of the FAE framework, we set our encoder as $\mathcal{E}:f\mapsto\tilde{\theta}$. If there exist multiple valid $\tilde{\theta}$ for $f\in F_{s}$, we can choose any one of them as the encoder’s output. As we see, the encoder $\mathcal{E}$ maps the function of the Barron space to a parameter set with bounded path norm. The decoder $\mathcal{D}:\theta\mapsto f({\mathbf{x}};\theta)$ is simply a parameter injection of a two-layer ReLU network. Hence, the autoencoder $\mathcal{D}\circ\mathcal{E}$ effectively serves as a two-layer ReLU network approximation. Next, we upper bound the density estimation loss $W_{1}(P,\widehat{P})$.
Theorem 3.5 (Density estimation over Barron space).
@@
The proof is provided in Appendix B.3. ∎
3.3 Methods

### v2

@@ Abstract
Recent advances in generative modeling—particularly diffusion models and flow matching—have achieved remarkable success in synthesizing discrete data such as images and videos. However, adapting these models to physical applications remains challenging, as the quantities of interest are continuous functions governed by complex physical laws. To address this, we introduce FunDiff, an efficient and robust framework for generative modeling in function spaces. FunDiff combines a latent diffusion process with a function autoencoder architecture to handle input functions with varying discretizations, generates continuous functions that can be evaluated at arbitrary locations, and seamlessly incorporate physical priors. These priors are enforced through architectural constraints or physics-informed loss functions, ensuring that generated samples satisfy fundamental physical laws. We theoretically establish minimax optimality guarantees for density estimation in function spaces, demonstrating that diffusion-based estimators achieve optimal convergence rates under suitable regularity conditions. We further demonstrate the practical effectiveness of FunDiff across diverse applications in fluid dynamics and solid mechanics. Empirical results indicate that our method can generate physically consistent samples with high fidelity to the target distribution, and exhibit robustness to noisy and low-resolution data.
Introduction
@@
Generative modeling, a cornerstone of modern machine learning, aims to learning and sampling complex data distributions $P(x)$ by constructing models that can generate new instances consistent with the learned distribution [1, 2]. Among recent advances in this domain, diffusion models and flow matching have emerged as powerful frameworks that achieve state-of-the-art performance. Diffusion models [3, 4] learn to reverse a forward noising process by gradually denoising data through a Markov chain, which iteratively transforms random noise into high-quality samples. In parallel, flow matching [5, 6] provides an alternative approach by learning continuous normalizing flows via optimal transport theory, directly modeling probability paths that transform a tractable prior distribution into the target data distribution.
@@
Diffusion models have shown remarkable success across a wide range of applications, including high-resolution image synthesis [7, 8, 9], audio generation [10], protein structure prediction [11], and molecular design [12]. Despite these advances, the focus has largely remained on discrete data types, such as images, text, and molecular graphs, whereas many scientific domains, especially in physics, demand the generation of continuous function data, where each sample represents a continuous function (e.g., a velocity or temperature field) defined over spatial and temporal domains. For instance, generating velocity fields in turbulent fluid dynamics, simulating deformation patterns in complex materials under external loadings, and forecasting temperature variations in climate models all require handling functions defined over continuous spatial and temporal domains. These problems inherently involve infinite-dimensional function spaces, a challenge that recent advances in operator learning frameworks have begun to address effectively [13, 14, 15, 16, 17, 18, 19]. This presents unique challenges:
@@
How can we extend generative modeling to generate continuous functions that can be evaluated at arbitrary points, while preserving important structural or physical properties inherent in scientific problems?
@@
Prior work in computer vision and graphics has extensively explored the concept of function-space generation through implicit neural representations and generative models, such as DeepSDF[20], NeRF[21], From Data to Functa[22], and SIREN[23], as well as their diffusion-based extensions including 3DShape2VecSet[24], Shap-E[25], and HyperDiffusion[26]. These methods treat signals such as shapes, radiance fields, or signed distance functions as continuous functions parameterized by neural networks, establishing an important foundation for generative modeling in continuous domains. Furthermore, recent works have attempted to extend the framework of generative models to function space, either by adapting existing algorithms to operate on discretized function data in $\mathbb{R}^{n}$ or by transforming input functions into a finite-dimensional latent space [27, 22, 28, 29, 30, 31]. Dupont et al. [32] pioneered this direction by treating data points as continuous functions and learning distributions over them using adversarial training. Multiple theoretical frameworks have emerged for function diffusion models, including Gaussian measures on Hilbert spaces [33], dimension-independent bounds for infinite-dimensional models [34], generalized diffusion processes in function spaces [35], and denoising diffusion operators with rigorous Gaussian processes and Langevin dynamics [36].Although these studies have demonstrated promising results for fixed discretizations, they generally struggle to scale to higher resolutions or to handle variable discretizations consistently, as those encountered in scientific computing, partial differential equation (PDE) modeling, or operator learning [14, 37, 38, 39, 40].
@@
In parallel, diffusion models have emerged as powerful tools for tackling inverse problems and modeling dynamical systems governed by PDEs. Recent works have developed methods for reconstructing spatiotemporal fields from sparse observations and partial supervision [41, 42, 43]. To address resolution limitations, several approaches have integrated neural operators with diffusion models [44, 45, 46], demonstrating improved performance in turbulence modeling and capture of multiscale phenomena. Furthermore, various strategies have attempted to incorporate physical priors directly into diffusion processes [47, 48, 49, 50, 42, 51, 52]. For instance, DiffusionPDE [42] introduces physics-based guidance during the posterior sampling stage, whereas PIDM [49] embeds PDE residuals directly into the training objective. However, enforcing PDE residuals is highly sensitive to solution accuracy—small numerical errors can destabilize both training and inference. Moreover, physical constraints are inherently defined in continuous space, but these diffusion models operate on discretized representations. This mismatch introduces discretization artifacts, often leading to inaccurate or ill-posed PDE residuals, especially in inverse problems where key state variables are only partially observed. These challenges underscore a fundamental gap between current data-driven generative models and the objective of producing physically consistent, continuous functions.
To address these challenges, we propose FunDiff, a novel framework for physics-informed generative modeling over function spaces. Our contributions are as follows. We introduce a unified architecture that combines latent diffusion processes with function autoencoders, enabling the high-quality data generation in a function-space representation. This design supports evaluation at arbitrary spatial locations while seamlessly incorporating physical priors through the function space representation. We establish theoretical guarantees for our approach by proving that diffusion-based models attain minimax-optimal rates for density estimation in infinite-dimensional spaces under mild smoothness assumptions. We demonstrate the practical effectiveness of FunDiff across diverse applications in fluid dynamics and solid mechanics. Our experiments show that the method generates physically consistent samples with high fidelity to target distributions while maintaining robustness to noisy and low-resolution input data. Taken together, our framework bridges the gap between theoretical guarantees in function spaces and practical neural network algorithms, establishing a principled approach to physics-informed generative modeling.
Results
In this section, we present theoretical results on generative models in function spaces, which establish the mathematical foundation of our work. Building on this theory, we introduce FunDiff, a practical framework for generating functions while incorporating physical priors. Then we demonstrate the effectiveness of our proposed FunDiff framework through a series of experiments on diverse physical systems. Specifically, we present four case studies with increasing complexity: (1) the Burgers’ equation, showcasing the incorporation of PDE constraints; (2) Kolmogorov flow, highlighting divergence-free field generation; (3) linear elasticity, illustrating symmetry preservation; and (4) turbulent mass transfer, demonstrating conditional generation capabilities. For notations, interested readers are encouraged to refer to the Appendix S1. It is important to emphasize that in all examples, FunDiff does not act as a neural PDE solver; rather, it learns to generate physically consistent continuous functions that approximate the distributions of solutions to selected PDE-governed systems when trained on such data.
Generative models over function spaces
@@ Results
In this section, we introduce a framework for generative modeling in function spaces. Given the inherently challenges of handling directly with infinitely-dimensional spaces, we begin by reducing the problem to a finite-dimensional setting, with preliminaries in Appendix S2. By embedding functions into a finite-dimensional vector space, the task of generating new functions can be reformulated as a well-studied vector generation task.
@@
where the expectation is taken over the randomness of the samples, and $\varepsilon(d,n)$ aligns with the known minimax estimation rates [53, 54]. For the continuous density class $\mathcal{C}({\mathbb{R}}^{d}):=\{\mu:{\mathbb{R}}^{d}\rightarrow[0,\infty)\penalty 10000\ :\penalty 10000\ \mu\penalty 10000\ \text{is continuous}\}$, we have a minimax density estimation rate
@@
[…]
@@
[…]
Density estimation loss decomposition
@@
[…]
@@
[…]
Illustrative example 1: Reproducing kernel Hilbert space (RKHS).
For a reproducing kernel Hilbert space (RKHS) $\mathcal{H}$, we denote its eigenvalues as $\mu_{1}\geqslant\mu_{2}\geqslant\ldots$, and their corresponding eigenfunctions as $\psi_{1},\psi_{2},\ldots$. For any function $f\in\mathcal{H}$, its eigen-decomposition has the form $f=\sum_{i=1}^{\infty}\theta_{i}\sqrt{\mu_{i}}\psi_{i}$ and its RKHS norm is defined as $\|f\|_{\mathcal{H}}^{2}=\sum_{i=1}^{\infty}\theta_{i}^{2}$. Define the RKHS unit ball as $\mathcal{B}:=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leqslant 1\}$.
@@
$f^{(k)}=\sum_{i=1}^{\infty}\theta_{i}^{(k)}\cdot\sqrt{\mu_{i}}\psi_{i},\penalty 10000\ \penalty 10000\ \penalty 10000\ (k=1,2,\ldots,n).$
@@
$\mathcal{E}:f\mapsto(\theta_{1},\theta_{2},\ldots,\theta_{D})^{\top},\penalty 10000\ \penalty 10000\ \penalty 10000\ \mathcal{D}:(\theta_{1},\theta_{2},\ldots,\theta_{D})^{\top}\mapsto\sum_{i=1}^{D}\theta_{i}\cdot\sqrt{\mu_{i}}\psi_{i}.$
[…]
@@
Theorem 1 (Density estimation over RKHS space).
Assume $P$ is a distribution supported on the RKHS unit ball $\mathcal{B}:=\{f\in\mathcal{H}:\|f\|_{\mathcal{H}}\leqslant 1\}$. By using the generative oracle $\mathcal{O}(D,n)$ and the function autoencoder (FAE) framework introduced above, the density estimation loss over RKHS space can be bounded as follows:
@@
The proof is provided in Appendix S3.2. ∎
Illustrative example 2: Barron space.
@@
Definition 2 (Path norm [56]).
@@
Definition 3 (Spectral norm).
@@
$f({\mathbf{x}})=\int_{{\mathbb{R}}^{d}}e^{i\langle{\mathbf{x}},\omega\rangle}\hat{F}(\omega)\,{\mathrm{d}}\omega,\penalty 10000\ \penalty 10000\ \forall{\mathbf{x}}\in\Omega.$
@@
Lemma 4.
[…]
@@
$\|{\mathbf{b}}_{k}\|_{1}=1,|c_{k}|\leqslant 1,\penalty 10000\ \penalty 10000\ \text{and}\penalty 10000\ \penalty 10000\ |a_{k}|\leqslant\frac{2\gamma(f)}{m}\leqslant\frac{2}{m}\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \text{holds for\penalty 10000\ \penalty 10000\ }\forall k\in[m].$
[…]
Theorem 5 (Density estimation over Barron space).
@@
The proof is provided in Appendix S3.3. ∎
Now we have proved the statistical upper bound of density estimation over the RKHS ball and Barron space. To further show the optimality of such an upper bound, we provide a statistical lower bound of density estimation over the RKHS ball, which matches its upper bound with only a negligible $\mathcal{O}(\mathrm{poly}(\log\log n))$ gap.
Theorem 6 (Minimax lower bound for Wasserstein estimation over an RKHS ball).
Assume $\mathcal{H}$ is an RKHS following $2\beta$-polynomial eigenvalue decay (Assumption 1)
and define the RKHS unit ball $\mathcal{B}:=\{f\in\mathcal{H}\penalty 10000\ :\penalty 10000\ \|f\|_{\mathcal{H}}\leqslant 1\}$. We observe $n$ i.i.d. samples $X_{1},X_{2},\ldots,X_{n}\sim P$, where the unknown distribution $P$ is supported on the RKHS ball $\mathcal{B}$. Let the statistical minimax rate
$\mathcal{R}_{n}:=\inf_{\widehat{P}_{n}}\sup_{P:\penalty 10000\ \mathrm{supp}(P)\subseteq\mathcal{B}}{\mathbb{E}}\left[W_{1}(P,\widehat{P}_{n})\right].$
Then there exists a constant $K(\beta,c,C)>0$, depending only on pure constants $\beta,c,C$ in the eigenvalue decay assumption, such that for all sufficiently large $n$,
$\mathcal{R}_{n}\geqslant K(\beta,c,C)\cdot\left(\frac{1}{\log n}\right)^{\beta}.$
Proof.
The proof is provided in Appendix S3.4. ∎
Notice that its corresponding statistical upper rate is $\left(\frac{\log\log n}{\log n}\right)^{\beta}$ according to our results in Theorem 1. It means that our upper and lower bounds only have a $\mathcal{O}(\mathrm{poly}(\log\log n))$ gap, which is negligible to our statistical rate.
FunDiff in practice
To realize the framework of function generative models outlined above, we propose FunDiff (Fig. 1), which consists of a novel FAE and a latent diffusion model to generate continuous functions. The FAE learns the mapping from discrete function representations to continuous ones, while also enabling the enforcement of prior physical constraints. The latent diffusion model is then trained directly on the latent manifold learned by the pretrained FAE encoder. Our generation process follows a simple two-step procedure: we first sample latent codes by integrating the reverse-time ODE from Gaussian noise to the target data distribution, and then transform these codes into continuous functions via the FAE decoder.
Physical prior integration.
$\mathcal{L}_{\text{PDE}}=|\mathcal{R}(\hat{f})|_{2}^{2},$
where $\mathcal{R}$ represents the PDE residual and $\hat{f}$ is the reconstructed function. The continuous representation of FAE enables automatic differentiation to compute all required derivatives, allowing constraint enforcement across the entire domain. It is worth noting that these physics-informed terms serve as regularizers rather than numerical solvers of the governing equations; they guide the learned FAE to produce output functions that are physically consistent.
Furthermore, some physical properties can be incorporated through architectural modifications. For periodic boundary conditions, one approach is to modify the coordinate embedding with Fourier bases of specific frequencies [59]. When handling divergence-free fields, particularly important in fluid dynamics applications, the FAE can be configured to learn a stream function $\psi$, from which velocity field components are derived as $u=\frac{\partial\psi}{\partial y},\quad v=-\frac{\partial\psi}{\partial x}$, thereby ensuring divergence-free flow. Symmetry constraints can be incorporated when needed through modifications to the decoder’s final layer.
Related work.
We emphasize that our approach fundamentally differs from existing physics-informed diffusion models [47, 48, 49, 50, 42, 51, 52], which impose physical constraints either by modifying the diffusion model’s training objective or by guiding the posterior sampling process. These methods operate on discretized outputs, where physical laws—typically enforced through PDE residuals—are subject to discretization errors. This not only leads to inaccurate constraint enforcement, but also introduces additional optimization challenges. Moreover, in inverse problems where key physical variables are unobserved, such residuals may be undefined at inference time, further limiting the applicability of these approaches. In contrast, our method enforces physical constraints within the FAE, which is fully decoupled from the training and inference processes of the latent diffusion model. Since the FAE provides a continuous neural representation, we can compute PDE residuals by automatic differentiation when needed. This design enables both exact and soft enforcement of constraints in continuous space, mitigating discretization-induced errors, and enabling more stable and physically consistent generation. Details on how to impose these constraints are provided in the following examples through comprehensive numerical experiments, demonstrating that our framework can generate diverse function samples while naturally preserving the underlying physical principles.